# Quantum Machine Learning: from NISQ to Fault Tolerance

### v1

@@ III Fault tolerant quantum computation (FTQC) algorithms
In the Fault-Tolerant Quantum Computation (FTQC) era, we envision a future where we have fully implemented quantum error correction. In this era, one of the most prominent and impactful applications is Shor’s algorithm Shor (1994). This algorithm has the remarkable ability to efficiently factorize large numbers, a capability that could potentially break the majority of the public-key cryptography systems currently used on the internet. Additionally, quantum computers, as originally envisioned by Richard Feynman Feynman et al. (1996), have the potential to significantly accelerate the simulation of quantum physics and chemistry.
@@
Despite these two remarkable applications, quantum computers have historically faced a problem of having limited use cases. While these two applications are groundbreaking, there is no concrete evidence that quantum computers will revolutionize our lives in the same way that classical computers have done Aaronson (2015).
@@
Recently, a series of newly proposed quantum algorithms have emerged, offering exciting possibilities for quantum computation. Among them, the Harrow-Hassidim-Lloyd (HHL) algorithm, introduced by its authors in 2008 Harrow et al. (2009), has gained significant attention. These novel algorithms not only hold the potential for significant speedups compared to classical counterparts but also show promise in addressing practical problems. These applications span various domains, including machine learning, clustering, classification, and the analysis of extensive datasets Aaronson (2015). Although, significant challenges are related to many algorithms designed for quantum machine learning, including dequantization and fast interfaces between classical and quantum processors.
@@
III.1 Quantum phase estimation and quantum principle component analysis
In the realm of quantum computing, quantum phase estimation is a pivotal quantum algorithm primarily designed to estimate the phase associated with an eigenvalue of a provided unitary operator Nielsen and Chuang (2010). The necessity to exponentiate the density matrix within quantum principle component analysis (PCA) aligns with the requirement for phase estimation.
In a formal context, suppose a unitary operator $U$ and its corresponding eigenvector |$\ket{\psi}~{}.$ The relationship is represented as: $U\ket{\psi}=e^{i2\pi\theta}\ket{\psi}~{}.$ Here, without loss of generality, we assume $0\leq\theta<1~{}.$ and the objective of Quantum Phase Estimation (QPE) is to accurately determine the value of $\theta$ while minimizing the number of operations required. The concept involves generating a binary approximation of $\theta$ through the application of quantum Fourier transformation. The preparatory steps are illustrated in FIG. 4.
@@ III.1 Quantum phase estimation and quantum principle component analysis
Beginning with |$\ket{\psi}$ and $n$ qubits initialized in the |$\ket{0}$ state: |$\ket{0}^{\otimes n}\otimes\ket{\psi}~{}.$ First one applies Hadamard gates to all |$\ket{0}$ qubits: $H^{\otimes n}\ket{0}^{\otimes n}=\frac{1}{2^{n/2}}\left(\ket{0}+\ket{1}\right)^{\otimes n}=\frac{1}{2^{n/2}}\sum_{k=0}^{2^{n-1}}\ket{k}~{}.$ One then decomposes $k~{}:$ $k=\sum_{j=0}^{n-1}k_{j}2^{j}~{}.$ The next step involves applying the unitary operator $U$ to |$\ket{\psi}$ while maintaining control over the other qubits. Practically, one applies $U^{2^{j}}$ on |$\ket{\psi}~{},$ only if $k_{j}=1~{},$ for $0\leq j\leq n-1~{}.$ Finally, this sequence of operations results in the state:
@@
The eigenstate |$\ket{\psi}$ is not needed at this stage. We proceed with the quantum Fourier transformation on the qubits |k::$\ket{k}~{}:$
$\begin{split}\frac{1}{2^{n/2}}\sum_{k=0}^{2^{n-1}}e^{i2\pi k\theta}\ket{k}=\frac{1}{2^{n}}\sum_{x=0}^{2^{n}-1}\sum_{k=0}^{2^{n-1}}\exp\left\{\frac{-i2\pi k}{2^{n}}\left(x-2^{n}\theta\right)\right\}\ket{x}~{}.\end{split}$ (11)
Now, one estimates $\theta$ by rounding $2^{n}\theta$ to the nearest integer. If $2^{n}\theta$ can be expressed as an integer, then one will obtain a precise result after performing measurements. On the other hand, If $\theta$ cannot be accurately expressed up to $n^{th}$ order binary, QPE will have an error. So the larger $n$ is, the smaller the error is. Because we have more space to approximate $\theta~{}.$ The good news is the error of QPE estimation is always bounded Nielsen and Chuang (2010). By approximating $\theta$, it’s possible to find the eigenvalue of the unitary operator. Consequently, since unitary operators can always be expressed as an exponential of a Hermitian operator, one can also determine the eigenvalue of the Hermitian matrix. This is the principle behind how quantum PCA calculates both the eigenvectors and eigenvalues of the density matrix.
One of the most fascinating applications in quantum machine learning is Quantum Principal Component Analysis. This method, akin to the popular Principal Component Analysis (PCA) used by companies like Netflix, enables the analysis of users’ preferences to offer tailored film recommendations. In a mathematical sense, the data provided by users is represented as vectors: $v_{j}~{},$ in a $d$ dimensional vector space. PCA’s objective is to identify the ideal axes for grouping this data into clusters. Another application could involve generating these vectors based on the fluctuations in stock prices for all stocks in the market, analyzing changes from discrete times $t_{i}$ to $t_{i+1}~{}$ Biamonte et al. (2017).
@@
Classically, this program is done as the following Biamonte et al. (2017). First, one calculate the covariance matrix: $\bm{C}=\sum_{j}\bm{v}_{j}\bm{v}_{j}^{T}~{},$ where $T$ represents the transverse matrix. In principle, one can understand the covariance matrix as a manifestation of the correlation between different component of the data. Now, one can diagonalize the covariance matrix and find its eigenvectors $c_{k}$ and corresponding eigenvalue $e_{k}$ : $\bm{C}=\sum_{k}e_{k}\bm{c}_{k}\bm{c}_{k}^{\dagger}~{}.$
@@
If only a few of the eigenvalues $c_{k}$ are large, and the remainder are small or zero, then the eigenvectors corresponding to those eigenvalues are called the principal components of $\bm{C}~{}.$ In more mathematical terms, if the diagonalized covariance matrix is sparse, then the non-zero eigenvalues are called the principle components. Each principal component represents an underlying common trend or form of correlation in the data, and decomposing a data vectors $\bm{v}$ in terms of principal components,
@@
If only a few of the eigenvalues $c_{k}$ are significantly large, and the rest are considerably small or zero, then the eigenvectors linked to these eigenvalues are termed the principal components of $\bm{C}$ In more technical terms, if the diagonalized covariance matrix is sparse, the non-zero eigenvalues represent the principal components. Each principal component characterizes an underlying common trend or form of correlation within the data. Decomposing a data vector $\bm{v}$ into these principal components through the equation: $\bm{v}=\sum_{k}\tilde{v}_{k}\bm{c}_{k}~{},$ enables compression of data representation and the anticipation of future behavior. Classical algorithms for PCA possess computational complexity and query complexity scaling as $\mathcal{O}(d^{2})~{}.$
@@
In the quantum version of this program, the initial step involves mapping these vectors $\bm{v}_{j}$ into quantum states: |$\ket{v_{j}}~{}.$ This mapping is facilitated by quantum random access memory (QRAM), a concept that will be further detailed in section III.4. A $d$ dimensional vector can be encoded in $\log{d}$ qubits. Assuming every vector $\bm{v}_{j}$ is equally probable or is randomly selected from a dataset containing $N$ vectors the resulting density matrix is: $\rho=\frac{1}{N}\sum_{j}|v_{j}\rangle\langle v_{j}|~{}.$ This density matrix essentially serves as the quantum equivalent of the covariance matrix $\bm{C}$ in the classical scenario. The subsequent step for quantum principal component analysis involves implementing two methods: density matrix exponentiation and quantum phase estimation. The former maps the density matrix $\rho$ into a unitary operator: $U_{\text{DME}}=e^{-i\theta\rho}~{},$ which enables us to utilize QPE to evaluate the eigenvalues and eigenstates.
@@
The latter method aims to find the eigenvectors and eigenvalues of the density matrix written in $U_{\text{DME}}$. By repeatedly sampling data and using the two aforementioned techniques, one can take any quantum version of a data vector |$\ket{\bm{v}_{j}}$ and decompose it into its principal components: |$\ket{v_{j}}\rightarrow\sum_{k}v_{k}\ket{c_{k}}~{}.$ It’s important to note that quantum principle component analysis scales as $\mathcal{O}(\log^{2}{d})$ in both computational complexity and query complexity, which represents a significant speedup compared to the classical version. In the study conducted in Huang et al. (2022a), the researchers present experimental evidence showcasing quantum advantage in PCA. However, a crucial prerequisite for this advantage is the availability of quantum data sourced from physical experiments, capable of generating density matrices. This contrasts with classical devices relying on QRAM for data processing. And the quantum advantage is measured in the number of queries to get the density matrix.
III.2 Classical algorithm framework for dequantizating QML models
@@
Despite the development of exciting quantum algorithms in QML, there are still open problems whether known QML algorithms provide new exponential speedups over classical algorithms for practically relevant instances of machine learning problems Ciliberto et al. (2018); Dunjko and Wittek (2020). This uncertainty is rooted in the challenge of identifying the source of this speedup. Some arguments suggest that QML’s exponential acceleration is solely attributed to its state preparation assumptions. Notably, the concept of “dequantization” has emerged, involving the development of classical algorithms that mimic the sampling assumptions of QML programs. Authors in this domain have created classical algorithms, often referred to as the dequantization of QML programs, which closely resemble their quantum counterparts but exhibit only polynomial differences in complexity. One notable example of dequantization involves the quantum recommendation system.
@@
The work by Kerenidis and Prakash on the quantum recommendation system Kerenidis and Prakash (2016) was noteworthy for addressing the caveats associated with the HHL algorithm, as pointed out by Scott Aaronson Aaronson (2015). They presented a comprehensive quantum algorithm directly comparable to classical algorithms. Initially, Kerenidis and Prakash’s quantum algorithm demonstrated exponential speedup compared to the best-known classical algorithms, but its provable exponential nature was uncertain. In 2019, Tang Tang (2019) introduced a classical algorithm that dequantized the quantum recommendation systems algorithm with only polynomially slower runtime, providing clarity on the provable nature of the speedup. These works highlighted the insight that the data structure fulfilling state preparation assumptions can also fulfill $\ell^{2}$-norm sampling assumptions (as defined in section 2.2 of Tang (2019)). More precisely, this is a $\log$-dimension time classical randomized linear algebra algorithms in the sample and query input model, which is an analogue of quantum state preparation assumptions coming from, for example, in most cases, QRAM). Consequently, a classical algorithm aiming to “match” the quantum algorithm can leverage these assumptions.
@@
In 2020, the authors of Chia et al. (2020) introduced an algorithmic framework for quantum-inspired classical algorithms focusing on close-to-low-rank matrices. This work generalized the series of results initiated by Tang’s quantum-inspired algorithm for recommendation systems Tang (2019). The motivation for these classical algorithms stemmed from quantum linear algebra algorithms and the quantum singular value transformation (SVT) framework Gilyén et al. (2019). The proposed classical algorithms for SVT exhibited runtime independent of input dimension under suitable quantum-inspired sampling assumptions. The results provided compelling evidence that, in the corresponding QRAM data structure input model, quantum SVT does not deliver exponential quantum speedups.
@@
In 2021, the authors of Tang (2021) introduced a new input model termed SQ access, representing a form of l2-norm sampling assumption. They developed a model to dequantize two influential QML algorithms, namely quantum principal component analysis Lloyd et al. (2014) and quantum supervised clustering Lloyd et al. (2013). In essence, they provided classical algorithms that, with classical SQ access assumptions replacing quantum state preparation assumptions, replicated the bounds and runtime of the corresponding quantum algorithms, albeit with a polynomial slowdown.
@@
On the other hand, in Huang et al. (2022a) the authors point out that the assumption in Tang (2021) does not work in the quantum physics environment setup, and it is possible that there are still quantum advantages with exponential separation. In Huang et al. (2022a), their quantum advantages are measured in terms of the number of queries to access the quantum state requiring principle component analysis. In Tang (2021), It presupposes the capability to retrieve any element of the exponentially large density matrix with exponentially precise accuracy, within a time frame that is polynomial. Achieving the capability already requires exponential numbers of queries to access the density matrix. See further detailed discussions in Cotler et al. (2021). In summary, the pursuit of identifying classical analogs for quantum machine learning applications presents a fascinating avenue for exploration. Investigating dequantization enables the identification of the constraints inherent in both quantum and classical machine learning, and aids in delineating the distinctions between quantum and classical algorithms.
III.3 Harrow-Hassidim-Lloyd algorithm

### v2

@@ 3 Fault tolerant quantum computation (FTQC) algorithms
In the Fault-Tolerant Quantum Computation (FTQC) era, we envision a future where we have fully implemented quantum error correction. In this era, one of the most prominent and impactful applications is Shor’s algorithm [117]. This algorithm has the remarkable ability to efficiently factorize large numbers, a capability that could potentially break the majority of the public-key cryptography systems currently used on the internet. Additionally, quantum computers, as originally envisioned by Richard Feynman [45], have the potential to significantly accelerate the simulation of quantum physics and chemistry.
@@
Despite these two remarkable applications, quantum computers have historically faced a problem of having limited use cases. While these two applications are groundbreaking, there is no concrete evidence that quantum computers will revolutionize our lives in the same way that classical computers have done [1].
@@
Recently, a series of newly proposed quantum algorithms have emerged, offering exciting possibilities for quantum computation. Among them, the Harrow-Hassidim-Lloyd (HHL) algorithm, introduced by its authors in 2008 [58], has gained significant attention. These novel algorithms not only hold the potential for significant speedups compared to classical counterparts but also show promise in addressing practical problems. These applications span various domains, including machine learning, clustering, classification, and the analysis of extensive datasets [1]. Although, significant challenges are related to many algorithms designed for quantum machine learning, including dequantization and fast interfaces between classical and quantum processors.
@@
3.1 Quantum phase estimation and quantum principle component analysis
In the realm of quantum computing, quantum phase estimation is a pivotal quantum algorithm primarily designed to estimate the phase associated with an eigenvalue of a provided unitary operator [102]. The necessity to exponentiate the density matrix within quantum principle component analysis (PCA) aligns with the requirement for phase estimation.
In a formal context, suppose a unitary operator $U$ and its corresponding eigenvector |$\ket{\psi}\leavevmode\nobreak\ .$ The relationship is represented as: $U\ket{\psi}=e^{i2\pi\theta}\ket{\psi}\leavevmode\nobreak\ .$ Here, without loss of generality, we assume $0\leq\theta<1\leavevmode\nobreak\ .$ and the objective of Quantum Phase Estimation (QPE) is to accurately determine the value of $\theta$ while minimizing the number of operations required. The concept involves generating a binary approximation of $\theta$ through the application of quantum Fourier transformation. The preparatory steps are illustrated in FIG. 4.
@@ 3.1 Quantum phase estimation and quantum principle component analysis
Beginning with |$\ket{\psi}$ and $n$ qubits initialized in the |$\ket{0}$ state: |$\ket{0}^{\otimes n}\otimes\ket{\psi}\leavevmode\nobreak\ .$ First one applies Hadamard gates to all |$\ket{0}$ qubits: $H^{\otimes n}\ket{0}^{\otimes n}=\frac{1}{2^{n/2}}\left(\ket{0}+\ket{1}\right)^{\otimes n}=\frac{1}{2^{n/2}}\sum_{k=0}^{2^{n-1}}\ket{k}\leavevmode\nobreak\ .$ One then decomposes $k\leavevmode\nobreak\ :$ $k=\sum_{j=0}^{n-1}k_{j}2^{j}\leavevmode\nobreak\ .$ The next step involves applying the unitary operator $U$ to |$\ket{\psi}$ while maintaining control over the other qubits. Practically, one applies $U^{2^{j}}$ on |$\ket{\psi}\leavevmode\nobreak\ ,$ only if $k_{j}=1\leavevmode\nobreak\ ,$ for $0\leq j\leq n-1\leavevmode\nobreak\ .$ Finally, this sequence of operations results in the state:
@@
The eigenstate |$\ket{\psi}$ is not needed at this stage. We proceed with the quantum Fourier transformation on the qubits |k::$\ket{k}\leavevmode\nobreak\ :$
$\begin{split}\frac{1}{2^{n/2}}\sum_{k=0}^{2^{n-1}}e^{i2\pi k\theta}\ket{k}=\frac{1}{2^{n}}\sum_{x=0}^{2^{n}-1}\sum_{k=0}^{2^{n-1}}\exp\left\{\frac{-i2\pi k}{2^{n}}\left(x-2^{n}\theta\right)\right\}\ket{x}\leavevmode\nobreak\ .\end{split}$ (11)
Now, one estimates $\theta$ by rounding $2^{n}\theta$ to the nearest integer. If $2^{n}\theta$ can be expressed as an integer, then one will obtain a precise result after performing measurements. On the other hand, If $\theta$ cannot be accurately expressed up to $n^{th}$ order binary, QPE will have an error. So the larger $n$ is, the smaller the error is. Because we have more space to approximate $\theta\leavevmode\nobreak\ .$ The good news is the error of QPE estimation is always bounded [102]. By approximating $\theta$, it’s possible to find the eigenvalue of the unitary operator. Consequently, since unitary operators can always be expressed as an exponential of a Hermitian operator, one can also determine the eigenvalue of the Hermitian matrix. This is the principle behind how quantum PCA calculates both the eigenvectors and eigenvalues of the density matrix.
One of the most fascinating applications in quantum machine learning is Quantum Principal Component Analysis. This method, akin to the popular Principal Component Analysis (PCA) used by companies like Netflix, enables the analysis of users’ preferences to offer tailored film recommendations. In a mathematical sense, the data provided by users is represented as vectors: $v_{j}\leavevmode\nobreak\ ,$ in a $d$ dimensional vector space. PCA’s objective is to identify the ideal axes for grouping this data into clusters. Another application could involve generating these vectors based on the fluctuations in stock prices for all stocks in the market, analyzing changes from discrete times $t_{i}$ to $t_{i+1}\leavevmode\nobreak$ [18].
@@
Classically, this program is done as the following [18]. First, one calculate the covariance matrix: $\bm{C}=\sum_{j}\bm{v}_{j}\bm{v}_{j}^{T}\leavevmode\nobreak\ ,$ where $T$ represents the transverse matrix. In principle, one can understand the covariance matrix as a manifestation of the correlation between different component of the data. Now, one can diagonalize the covariance matrix and find its eigenvectors $c_{k}$ and corresponding eigenvalue $e_{k}$ : $\bm{C}=\sum_{k}e_{k}\bm{c}_{k}\bm{c}_{k}^{\dagger}\leavevmode\nobreak\ .$
@@
If only a few of the eigenvalues $c_{k}$ are large, and the remainder are small or zero, then the eigenvectors corresponding to those eigenvalues are called the principal components of $\bm{C}\leavevmode\nobreak\ .$ In more mathematical terms, if the diagonalized covariance matrix is sparse, then the non-zero eigenvalues are called the principle components. Each principal component represents an underlying common trend or form of correlation in the data, and decomposing a data vectors $\bm{v}$ in terms of principal components,
@@
If only a few of the eigenvalues $c_{k}$ are significantly large, and the rest are considerably small or zero, then the eigenvectors linked to these eigenvalues are termed the principal components of $\bm{C}$ In more technical terms, if the diagonalized covariance matrix is sparse, the non-zero eigenvalues represent the principal components. Each principal component characterizes an underlying common trend or form of correlation within the data. Decomposing a data vector $\bm{v}$ into these principal components through the equation: $\bm{v}=\sum_{k}\tilde{v}_{k}\bm{c}_{k}\leavevmode\nobreak\ ,$ enables compression of data representation and the anticipation of future behavior. Classical algorithms for PCA possess computational complexity and query complexity scaling as $\mathcal{O}(d^{2})\leavevmode\nobreak\ .$
@@
In the quantum version of this program, the initial step involves mapping these vectors $\bm{v}_{j}$ into quantum states: |$\ket{v_{j}}\leavevmode\nobreak\ .$ This mapping is facilitated by quantum random access memory (QRAM), a concept that will be further detailed in section 3.4. A $d$ dimensional vector can be encoded in $\log{d}$ qubits. Assuming every vector $\bm{v}_{j}$ is equally probable or is randomly selected from a dataset containing $N$ vectors the resulting density matrix is: $\rho=\frac{1}{N}\sum_{j}|v_{j}\rangle\langle v_{j}|\leavevmode\nobreak\ .$ This density matrix essentially serves as the quantum equivalent of the covariance matrix $\bm{C}$ in the classical scenario. The subsequent step for quantum principal component analysis involves implementing two methods: density matrix exponentiation and quantum phase estimation. The former maps the density matrix $\rho$ into a unitary operator: $U_{\text{DME}}=e^{-i\theta\rho}\leavevmode\nobreak\ ,$ which enables us to utilize QPE to evaluate the eigenvalues and eigenstates.
@@
The latter method aims to find the eigenvectors and eigenvalues of the density matrix written in $U_{\text{DME}}$. By repeatedly sampling data and using the two aforementioned techniques, one can take any quantum version of a data vector |$\ket{\bm{v}_{j}}$ and decompose it into its principal components: |$\ket{v_{j}}\rightarrow\sum_{k}v_{k}\ket{c_{k}}\leavevmode\nobreak\ .$ It’s important to note that quantum principle component analysis scales as $\mathcal{O}(\log^{2}{d})$ in both computational complexity and query complexity, which represents a significant speedup compared to the classical version. In the study conducted in [67], the researchers present experimental evidence showcasing quantum advantage in PCA. However, a crucial prerequisite for this advantage is the availability of quantum data sourced from physical experiments, capable of generating density matrices. This contrasts with classical devices relying on QRAM for data processing. And the quantum advantage is measured in the number of queries to get the density matrix.
3.2 Classical algorithm framework for dequantizating QML models
@@
Despite the development of exciting quantum algorithms in QML, there are still open problems whether known QML algorithms provide new exponential speedups over classical algorithms for practically relevant instances of machine learning problems [34, 40]. This uncertainty is rooted in the challenge of identifying the source of this speedup. Some arguments suggest that QML’s exponential acceleration is solely attributed to its state preparation assumptions. Notably, the concept of “dequantization” has emerged, involving the development of classical algorithms that mimic the sampling assumptions of QML programs. Authors in this domain have created classical algorithms, often referred to as the dequantization of QML programs, which closely resemble their quantum counterparts but exhibit only polynomial differences in complexity. One notable example of dequantization involves the quantum recommendation system.
@@
The work by Kerenidis and Prakash on the quantum recommendation system [72] was noteworthy for addressing the caveats associated with the HHL algorithm, as pointed out by Scott Aaronson [1]. They presented a comprehensive quantum algorithm directly comparable to classical algorithms. Initially, Kerenidis and Prakash’s quantum algorithm demonstrated exponential speedup compared to the best-known classical algorithms, but its provable exponential nature was uncertain. In 2019, Tang [121] introduced a classical algorithm that dequantized the quantum recommendation systems algorithm with only polynomially slower runtime, providing clarity on the provable nature of the speedup. These works highlighted the insight that the data structure fulfilling state preparation assumptions can also fulfill $\ell^{2}$-norm sampling assumptions (as defined in section 2.2 of [121]). More precisely, this is a $\log$-dimension time classical randomized linear algebra algorithms in the sample and query input model, which is an analogue of quantum state preparation assumptions coming from, for example, in most cases, QRAM). Consequently, a classical algorithm aiming to “match” the quantum algorithm can leverage these assumptions.
@@
In 2020, the authors of [32] introduced an algorithmic framework for quantum-inspired classical algorithms focusing on close-to-low-rank matrices. This work generalized the series of results initiated by Tang’s quantum-inspired algorithm for recommendation systems [121]. The motivation for these classical algorithms stemmed from quantum linear algebra algorithms and the quantum singular value transformation (SVT) framework [47]. The proposed classical algorithms for SVT exhibited runtime independent of input dimension under suitable quantum-inspired sampling assumptions. The results provided compelling evidence that, in the corresponding QRAM data structure input model, quantum SVT does not deliver exponential quantum speedups.
@@
In 2021, the authors of [122] introduced a new input model termed SQ access, representing a form of l2-norm sampling assumption. They developed a model to dequantize two influential QML algorithms, namely quantum principal component analysis [93] and quantum supervised clustering [92]. In essence, they provided classical algorithms that, with classical SQ access assumptions replacing quantum state preparation assumptions, replicated the bounds and runtime of the corresponding quantum algorithms, albeit with a polynomial slowdown.
@@
On the other hand, in [67] the authors point out that the assumption in [122] does not work in the quantum physics environment setup, and it is possible that there are still quantum advantages with exponential separation. In [67], their quantum advantages are measured in terms of the number of queries to access the quantum state requiring principle component analysis. In [122], It presupposes the capability to retrieve any element of the exponentially large density matrix with exponentially precise accuracy, within a time frame that is polynomial. Achieving the capability already requires exponential numbers of queries to access the density matrix. See further detailed discussions in [37]. In summary, the pursuit of identifying classical analogs for quantum machine learning applications presents a fascinating avenue for exploration. Investigating dequantization enables the identification of the constraints inherent in both quantum and classical machine learning, and aids in delineating the distinctions between quantum and classical algorithms.
3.3 Harrow-Hassidim-Lloyd algorithm